# Forecasting CPI inflation under economic policy and geo-political uncertainties

### v1

@@ 5.5 Conformalized Prediction Intervals
Next, we focus on quantifying the uncertainty associated with the forecast generated by the FEWNet approach using a conformal prediction interval. The non-parametric conformal prediction approach, introduced in [66], is a set of methods that usually take an uncertainty score and turn it into a probabilistic band, including the true outcome. In other words, this framework converts point estimations into a prediction region. One of the significant advantages of conformal prediction intervals is that this model-agnostic procedure guarantees coverage. In the time series setup, this method generates the prediction interval by leveraging the sequential ordering of the data. Given the training set $\left\{Y_{t},\tilde{X}_{t}\right\}_{t=1}^{N}$ where $Y_{t}$ is the target series, and $\tilde{X}_{t}$ is the set of features, including lagged historical values of $Y_{t}$ and the covariates, we fit the FEWNet framework and an uncertainty model $\widehat{\Xi}$ on $\tilde{X}_{t}$ to generate a scaler notion of uncertainty. The conformal score $\mathcal{S}_{t}$ can be calculated as:
$\mathcal{S}_{t}=\frac{\left|Y_{t}-\operatorname{FEWNet}\left(\tilde{X}_{t}\right)\right|}{\widehat{\Xi}(\tilde{X}_{t})}$
@@
^$\widehat{\mathbb{Q}}_{t}=inf\left\{q:\ \frac{1}{\min\left(\kappa,t^{{}^{\prime}}-1\right)+1}\sum_{t^{{}^{\prime}}-1}^{t-1}{\mathcal{S}_{t^{{}^{\prime}}}\mathbb{1}\left(t^{{}^{\prime}}\geq t-\ \kappa\right)\geq 1-\alpha}\right\}$
@@
$\mathbb{C}\left(\tilde{X}_{t}\right)=\left[\operatorname{FEWNet}\left(\tilde{X}_{t}\right)-{\widehat{\mathbb{Q}}}_{t}\widehat{\Xi}\left(\tilde{X}_{t}\right),\operatorname{FEWNet}\left(\tilde{X}_{t}\right)+{\widehat{\mathbb{Q}}}_{t}\ \widehat{\Xi}\left(\tilde{X}_{t}\right)\right].$
@@
In this section, we present the conformal prediction intervals for the 24-month ahead CPI inflation forecasts generated by the FEWNet model using the validation data points (2019-12 to 2021-11). The conformal prediction interval of the FEWNet framework, calculated based on the forecast estimates and residuals for the validation set, is demonstrated in Figure 5 along with the point estimate of the forecast produced by FEWNet and the best-performing baseline models for the BRIC countries. As observed in the figure the prediction intervals vary in width across different datasets. For instance, in the case of the 24-month forecast for Brazil, the width of the interval averages around 8.18, for Russia, it hovers around 7.93, for India, it reaches the minimum of 3.88; and for China, it averages around 7.60. A similar analysis of the conformal prediction intervals for the semi-long term forecast horizon is provided in Appendix 7.3. The overall analysis thus demonstrates a form of uncertainty quantification for the forecast of CPI inflation series for BRIC countries and shows the structural shift in the measurement of prediction uncertainty across different countries.
@@
Figure 5: The plot shows the ground truth CPI inflation data (red dots), forecast generated by the best (yellow) and second-best (orange) baseline models for each country, conformal prediction interval (green shaded), and forecast (blue) generated by the FEWNet model for the long-term forecast horizon in BRIC Countries
@@ 6 Conclusion and Discussion
Historically, emerging economies have exhibited higher and more enduring inflation rates when compared to developed countries. Most emerging economies have implemented an inflation-targeting monetary policy framework in which inflation serves as the intermediate target. This paper outlines a novel machine learning framework, namely FEWNet, that combines the MODWT process with the ARNNx model for long-term CPI inflation forecasting for the BRIC countries under macroeconomic and geopolitical uncertainties. The proposed method demonstrates consistent performances for 3 out of 4 major BRIC economies for both the semi-long run and long-run forecast horizons. Since the CPI inflation data of China presents a stationary behavior, our proposal failed to outperform benchmark forecasters that work exceptionally well on stationary time series. Moreover, for Brazil, Russia, and India it produces the most accurate performance, outperforming the classical and SOTA machine learning or deep learning algorithms. We also present a framework for time series feature engineering using different economic filters (HP and CF band-pass filters) that can be very easily leveraged in the proposed architecture, which supports the inclusion of exogenous regressors within its flexible design. A combination of the multiresolution analysis and autoregressive neural networks are used to predict nonlinearity, nonstationarity, and non-gaussian time series of inflation. MODWT decomposition produces new time series that contains system and slow dynamics that easily separate signal from noise. Each time series is further modeled with ARNNx. Theoretically, we show that under certain conditions, the proposed FEWNet reduces the empirical risk, resulting in improved forecast accuracy.
@@
Although FEWNet is designed to predict CPI inflation of BRIC countries, it can be seamlessly extended for developed countries and can include other uncertainty variables like Twitter-based uncertainties, financial stress indicators, world uncertainty index, etc. This framework can also be leveraged for forecasting other macroeconomic variables like GDP, gold price, unemployment rate, and others. The FEWNet framework is more effective for generating forecasts for non-stationary and non-linear time series data that are evident in macroeconomics; therefore, it can potentially serve as a viable alternative to forecasting strategy for central banks. In light of our results, further analysis of the FEWNet’s performance in predicting other macroeconomic time series as well as inflation in a multivariate forecasting setting seems to be a promising avenue for future research.

### v2

@@ 5.7 Conformalized Prediction Intervals
Next, we focus on quantifying the uncertainty associated with the forecast generated by the FEWNet approach using a conformal prediction interval. The non-parametric conformal prediction approach, introduced in Vovk et al. (2005), is a set of methods that usually take an uncertainty score and turn it into a probabilistic band, including the true outcome. In other words, this framework converts point estimations into a prediction region. One of the significant advantages of conformal prediction intervals is that this model-agnostic procedure guarantees coverage. In the time series setup, this method generates the prediction interval by leveraging the sequential ordering of the data. Given the training set $\left\{Y_{t},\tilde{X}_{t}\right\}_{t=1}^{N}$ where $Y_{t}$ is the target series, and $\tilde{X}_{t}$ is the set of features, including lagged historical values of $Y_{t}$ and the covariates, we fit the FEWNet framework and an uncertainty model $\widehat{\Xi}$ on $\tilde{X}_{t}$ to generate a scaler notion of uncertainty. The conformal score $\mathcal{S}_{t}$ can be calculated as:
$\mathcal{S}_{t}=\frac{\left|Y_{t}-\operatorname{FEWNet}\left(\tilde{X}_{t}\right)\right|}{\widehat{\Xi}(\tilde{X}_{t})}$ (12)
@@
^$\widehat{\mathbb{Q}}_{t}=inf\left\{q:\ \frac{1}{\min\left(\kappa,t^{{}^{\prime}}-1\right)+1}\sum_{t^{{}^{\prime}}=1}^{t-1}{\mathcal{S}_{t^{{}^{\prime}}}\mathbb{1}\left(t^{{}^{\prime}}\geq t-\ \kappa\right)\geq 1-\alpha}\right\}$ (13)
@@
$\mathbb{C}\left(\tilde{X}_{t}\right)=\left[\operatorname{FEWNet}\left(\tilde{X}_{t}\right)-{\widehat{\mathbb{Q}}}_{t}\widehat{\Xi}\left(\tilde{X}_{t}\right),\operatorname{FEWNet}\left(\tilde{X}_{t}\right)+{\widehat{\mathbb{Q}}}_{t}\ \widehat{\Xi}\left(\tilde{X}_{t}\right)\right].$ (14)
@@
In this section, we present the conformal prediction intervals for the 24 months ahead of CPI inflation forecasts generated by the FEWNet model using the validation data points (2019-12 to 2021-11). The conformal prediction interval of the FEWNet framework, calculated based on the forecast estimates and residuals for the validation set, is demonstrated in Figure 7 along with the point estimate of the forecast produced by FEWNet and the best-performing baseline models for the BRIC countries. Although the AR model ranked $2^{nd}$ based on the performance metrics however, it significantly fails to capture the movements in the test CPI inflation series and may not be an effective forecasting strategy. As observed in the figure, the prediction intervals vary in width across different datasets. For instance, in the case of the 24-month forecast for Brazil, the width of the interval averages around 8.18; for Russia, it hovers around 7.93; for India, it reaches the minimum of 3.88; and for China, it averages around 7.60. A similar analysis of the conformal prediction intervals for the semi-long-term forecast horizon is provided in Appendix 8.5. The overall analysis thus demonstrates a form of uncertainty quantification for the forecast of CPI inflation series for BRIC countries and shows the structural shift in the measurement of prediction uncertainty across different countries.
@@
Figure 7: The plot shows the ground truth CPI inflation data (red dots), 24 months point forecasts generated by FEWNet (blue line), AR (green line), XGBoost (purple line), and the conformal prediction interval of FEWNet (yellow shaded) for BRIC Countries.
6 Policy Implications
Most of the BRIC countries, except China, have adopted inflation targeting as a framework for monetary policy, in which inflation serves as an intermediate target. Accurately forecasting of inflation in such countries is essential for maintaining the credibility of these targets. By reliably forecasting inflation, central bankers in these countries can communicate their policy intentions, manage public expectations, and build trust in their commitment to price stability. However, as can be seen from Tables 2 and 3, inflation trends in BRIC countries are characterized by complex, volatile, and non-linear patterns influenced by numerous internal and external factors. Two major factors that lead to such patterns are EPU and GPRC, which affect the demand or supply side of the economy Adeosun et al. (2023), leading to large fluctuations in inflation Balcilar et al. (2017). For example, BRIC countries depend significantly on commodities such as oil, gas, and agricultural products, among many others. Consequently, fluctuations in global commodity prices due to geopolitical risk can directly impact their inflation rates. Additionally, EPU can cause increased volatility in financial markets, affecting exchange rates and commodity prices. For BRIC countries, which are often significant commodity exporters, this can lead to fluctuating import and export prices, impacting inflation. Moreover, geopolitical tensions (e.g., the Russia-Ukraine War) can lead to economic sanctions, trade disruptions, and shifts in global economic alliances, further affecting inflation through these channels. From the central banks’ point of view, high EPU and GPRC make it harder for central banks to set appropriate interest rates. Reactive monetary policies may become less effective, leading to more pronounced inflation volatility and making it harder to maintain price stability Balcilar et al. (2017). Given the dynamic relation between inflation, EPU, and GPRC in BRIC countries, it is essential to incorporate these uncertainty measures as exogenous variables in inflation forecasting models. Ignoring them can make the forecasting process even more challenging and result in inaccuracies for central banks Binder & Sekkel (2024); Nasir (2020). Traditional forecasting models often struggle to adapt to sudden changes and capture the non-linear dynamics characteristic of these environments. In contrast, machine learning models, particularly ensemble models, offer significant advantages for inflation forecasting in uncertain times due to their ability to process vast amounts of data and identify complex patterns John et al. (2020); Medeiros et al. (2021).
In this paper, by employing the wavelet coherence analysis, we show a significant causal relation between inflation, EPU, and GPRC in BRIC countries. Excluding them in the inflation forecasting models will lead to model misspecification, resulting in incorrect forecasts [^1] For instance, during both the global financial crisis and the COVID-19 pandemic, central banks in advanced economies experienced significant forecasting errors, notably underestimating inflation Alessi et al. (2014). Researchers at the European Central Bank identified a major cause for these errors: the neglect of various financial and uncertainty indicators and the failure to account for non-linear dynamics in forecasting models Kenny & Morgan (2011).. Because the transmission channels of inflation and monetary policy during periods of high EPU (such as the global financial crisis 2007-09) and GPRC change substantially Benchimol (2016). This argument is further supported by the findings of Benchimol & El-Shagi (2020) where they conclude that while predicting inflation for a country like Israel, the policymakers should consider the current level of terrorism (one of the aspects of GPRC) in their inflation forecasting model.
Furthermore, we show that our proposed model, FEWNet, performed superior in forecasting CPI inflation for Brazil, Russia, and India for semi-long-run and long-run horizons. FEWNet integrates economic filters like HP and CF with the wavelet decomposed autoregressive neural networks to significantly improve the forecast accuracy compared to traditional models. Specifically, we can conclude that FEWNet is more suitable for countries where inflation is characterized by non-stationarity and non-linearity due to heightened volatility and other complex dynamics. Given these findings, we recommend that when policymakers at central banks in BRIC countries evaluate the future path of inflation, they should include uncertainty and risk variables like EPU and GPRC and adopt advanced machine learning models like the FEWNet model for more accurate inflation forecasting. This is also evident from the theory (Section 4) and empirical study (Section 5.5) of the empirical risk for the proposed FEWNet model. With more accurate inflation forecasts, central banks in these countries can better align their monetary policies with actual economic conditions. This reduces the risk of policy errors, such as setting interest rates that are too high or too low relative to the economic context. Advanced models like FEWNet enable central banks to more effectively anticipate and respond to inflationary trends. This proactive approach facilitates timely and targeted interventions, especially in countries with complex and dynamic economic conditions, helping to stabilize inflation expectations and maintain economic stability.
@@ 7 Conclusion and Discussion
This paper outlines a novel machine learning framework, namely FEWNet, that combines the MODWT process with the ARNNx model for long-term CPI inflation forecasting for the BRIC countries under macroeconomic and geopolitical uncertainties. The proposed method demonstrates consistent performances for most of the BRIC economies for both the semi-long-run and long-run forecast horizons. Since the CPI inflation data of China presents a stationary behavior, our proposal failed to outperform all the benchmark forecasters that work exceptionally well on stationary time series. Moreover, it produces the most accurate performance for Brazil, Russia, and India, outperforming state-of-the-art machine learning or deep learning algorithms. We also present a framework for time series feature engineering using different economic filters (HP and CF band-pass filters) that have been leveraged in the proposed architecture, which supports the inclusion of exogenous regressors within its flexible design. A combination of the multiresolution analysis and autoregressive neural networks are used to predict non-linearity, non-stationarity, and non-gaussian time series of inflation. MODWT decomposition produces a trend time series and multiple detail series that contain system and slow dynamics, which easily separate signal from noise. Each time series is further modeled with ARNNx. Both theoretically and empirically, we show that under certain conditions, the proposed FEWNet reduces the empirical risk, resulting in improved forecast accuracy.
@@
While FEWNet is designed here to predict CPI inflation in BRIC countries, its flexible framework allows for seamless adaptation to other developed countries, including the OECD countries, and the inclusion of various additional uncertainty variables like Twitter (X)-based uncertainty measures, financial stress indicators, and broader indices like the World Uncertainty Index. This can be considered future research for this study. Moreover, the FEWNet framework is not limited to inflation forecasting alone. It holds significant potential for forecasting a wide array of macroeconomic variables like Gross Domestic Product (GDP), gold prices, unemployment rates, and more. Its strength lies in effectively handling non-stationary and non-linear time series, which are common in macroeconomic contexts. Given its capabilities, FEWNet presents itself as a robust alternative for central banks and policymakers, which often require accurate forecasts in complex economic environments. In light of our results, further analysis of the FEWNet’s performance in predicting other macroeconomic time series as well as inflation in a multivariate forecasting setting seems to be a promising avenue for future research.
Acknowledgement
The authors would like to acknowledge the Associate Editor and two experienced reviewers for insightful suggestions that improved the paper. The authors want to thank Madhurima Panja of IIIT Bangalore, India, for helping with the graphics presented in this manuscript.
8 Appendix
8.1 Discrete Wavelet Transformation:
The FEWNet framework leverages a form of discrete wavelet transformation (DWT) to first denoise the CPI inflation series for the BRIC countries, followed by modeling of the resulting components with the auxiliary information (trend and cycle generated through HP and CF filters) using ARNNx Faraway & Chatfield (1998) in an ensemble setup. In particular, we focus on the ‘maximal overlapping’ version of the DWT, popularly known as MODWT. In the literature, the DWT framework has been widely used in various applications for compressing digital images, smoothing signals Percival & Walden (2000); Walden (2001), material science Li et al. (2020), atmosphere Percival & Mofjeld (1997), energy Yang & Wang (2021), economics, and geophysics Zhu et al. (2014) among many others. We begin with a brief description of the DWT approach that forms the mathematical basis of the MODWT to be used in our proposed FEWNet framework.
The DWT approach defines a class of orthogonal wavelet transformation Percival & Guttorp (1994). Let $\{h_{l};l=0,1,\ldots,L-1\}$ denote a finite length discrete high-pass (wavelet) filter such that it satisfies the unit energy assumptions and it sums up to zero, i.e.,
$\sum_{l=0}^{L-1}h_{l}^{2}=1\;\text{and}\;\sum_{l=0}^{L-1}h_{l}=0.$ (15)
Furthermore, the wavelet filters are also restricted to be orthogonal to even shifts. Thus mathematically we can denote this as
$\sum_{l=0}^{L-1}h_{l}h_{l+2\tilde{n}}=0\;\;\text{for any non zero integer }\tilde{n}.$ (16)
Alongside the high-pass filter, having a low-pass (scaling) filter $\{g_{l};l=0,1,\ldots,L-1\}$ is also crucial for decomposing an observed time series into high-frequency oscillations and low-frequency details. These scaling filters also satisfy the unit energy assumptions and are orthogonal to even shifts (as in Eqs. (15), (16)). Furthermore, for all the wavelet filters, the low-pass filter coefficients are determined by the ‘quadrature mirror’ relationship, i.e., $g_{l}=(-1)^{l+1}h_{L-1-l}\;\;\text{or}\;\;h_{l}=(-1)^{l}g_{L-1-l};\;l=0,1,\ldots,L-1.$
For the construction of DWT coefficients from a given series, the use of the “pyramid algorithm” is prevalent Percival & Guttorp (1994); Percival & Walden (2000). Suppose we denote the economic time series to be transformed as $Y=\{Y_{t},t=1,2,\ldots\mathcal{N}\}$ of length $\left(\mathcal{N}=2^{\mathcal{K}}\right)$. In the pyramid algorithm, given the series $Y$, the wavelet filters $h_{l}$, and the scaling filters $g_{l}$, the first iteration performs a down-sampling operation and convolutes the data vector with each of the filters to obtain the resulting wavelet and scaling coefficient. In the subsequent step, similar down-sampling and convolution take place on the scaling coefficients of the previous step. Following this procedure iteratively the $\tilde{k}^{th}$ stage output of the pyramid algorithm can be expressed as
$\varpi_{\tilde{k},t}=\sum_{l=0}^{L_{\tilde{k}}-1}h_{\tilde{k},l}Y_{2^{\tilde{k}}(t+1)-l\text{ mod }\mathcal{N}}\;\;\text{and}\;\;v_{\tilde{k},t}=\sum_{l=0}^{L_{\tilde{k}}-1}g_{\tilde{k},l}Y_{2^{\tilde{k}}(t+1)-l\text{ mod }\mathcal{N}},$ (17)
where $\varpi_{\tilde{k},t}$ and $v_{\tilde{k},t}$ are the $\tilde{k}^{th}\;(\tilde{k}=1,2,\ldots,\mathcal{K})$ level wavelet and scaling coefficients respectively. This procedure can be repeated up to $\operatorname{log}_{2}(\mathcal{N})$ times. Thus, using the DWT on the given time series, we can represent the discrete wavelet coefficients in matrix notation as
$\varpi=\mathcal{W}Y,$ (18)
where $\mathcal{W}$, an orthogonal matrix of order $\mathcal{N}\times\mathcal{N}$, comprises of wavelet and scaling filters arranged in a row-by-row basis. The wavelet coefficients $\varpi$ can be arranged into $\mathcal{K}+1$ vectors as $\varpi=[\varpi_{1},\varpi_{2},\ldots,\varpi_{\mathcal{K}},v_{\mathcal{K}}]^{T}$ with each of $\varpi_{\tilde{k}},\;({\tilde{k}}=1,2,\ldots,\mathcal{K})$, a vector of length $\mathcal{N}/2^{\tilde{k}}$, consists of the wavelet coefficients associated with changes on a scale of length $2^{\tilde{k}-1}$ and the vector $v_{\mathcal{K}}$ of length $\mathcal{N}/2^{\mathcal{K}}$ is associated with averages on a scale of length $2^{\mathcal{K}-1}$.
Although the DWT is a useful mathematical tool for decomposing discrete time series into simpler parts using high-pass and low-pass filters, it suffers from certain serious limitations. To begin with, the DWT approach down-samples the signal at each iteration, and it can be repeated for a maximum of $\operatorname{log}_{2}\left(\mathcal{N}\right)$, thus restricting the size of the signal to be an exact power of 2. Moreover, the wavelet and scaling coefficients generated from a DWT convoluted signal do not scale and are not shift invariant. This restricts the universal application of the DWT algorithm for arbitrary time series. To overcome these deficiencies of the DWT framework, a modified maximal overlapping version of DWT was proposed Percival & Guttorp (1994); Percival & Walden (2000); Walden (2001).
8.2 MODWT approach
The maximal overlapping discrete wavelet transform (MODWT) is an improved, non-subsampled version of the DWT framework. The MODWT overcomes the limitations of the DWT approach and is invariant to circular shifts. This non-decimated wavelet transform is suitable for handling the non-stationary behavior of a discrete-time process and is highly redundant. Moreover, the variance estimator pertaining to the MODWT process is asymptotically more efficient than that of the DWT Percival & Guttorp (1994). Thus, in our study, we utilize the transition and time-invariant Pesquet et al. (1996) MODWT transformation for denoising the CPI inflation series. The mathematical formulation of the MODWT algorithm can be extended from the DWT approach and is described below.
For the MODWT algorithm, we define the rescaled version of filters as $\tilde{h}_{l}=h_{l}/2^{l}$ and $\tilde{g}_{l}=g_{l}/2^{l}$ and employ the pyramid algorithm with the time series $Y$, wavelet filter $\tilde{h}_{l}$, and scaling filter $\tilde{g}_{l}$. Similar to the DWT approach in the non-decimated MODWT framework, the pyramid algorithm convolutes the data with the re-normalized filters and generates the resulting coefficients of length $\mathcal{N}$ as
$\tilde{\varpi}_{\tilde{k},t}=\sum_{l=0}^{L_{\tilde{k}}-1}\tilde{h}_{\tilde{k},l}Y_{(t+1)-l\text{ mod }\mathcal{N}}\;\;\text{and}\;\;\tilde{v}_{\tilde{k},t}=\sum_{l=0}^{L_{\tilde{k}}-1}\tilde{g}_{\tilde{k},l}Y_{(t+1)-l\text{ mod }\mathcal{N}},$ (19)
where $L_{\tilde{k}}=(2^{\tilde{k}}-1)(L-1)+1$. The wavelet ($\tilde{\varpi}_{\tilde{k},t}$) and scaling ($\tilde{v}_{\tilde{k},t}$) coefficients obtained by decomposing the data vector $Y$ using the MODWT approach into $\mathcal{K}$ levels can be formulated as
$\tilde{\varpi}=\tilde{\mathcal{W}}Y$ (20)
where the orthonormal matrix $\tilde{\mathcal{W}}=[\tilde{\mathcal{W}}_{1},\tilde{\mathcal{W}}_{2},\ldots,\tilde{\mathcal{W}}_{\mathcal{K}},\tilde{V}_{\mathcal{K}}]^{T}$ is made up of $\mathcal{K}+1$ submatrices of size $\mathcal{N}\times\mathcal{N}$ with $\tilde{\mathcal{W}}_{\tilde{k}},\;(\tilde{k}=1,2,\ldots,\mathcal{K})$ and $\tilde{V}_{\mathcal{K}}$ denoting circularly shifted wavelet and scaling filters respectively. Furthermore, the wavelet and scaling coefficients can be arranged in $\mathcal{K}+1$ vectors as $\tilde{\varpi}=[\tilde{\varpi}_{1},\tilde{\varpi}_{2},\ldots,\tilde{\varpi}_{\mathcal{K}},\tilde{v}_{\mathcal{K}}]^{T}$ in a similar manner as in Eq. (18) of DWT approach.
Following the MODWT decomposition, an additive multi-resolution analysis (MRA) of the observed time series $Y$ can be formulated as
$Y_{t}=\sum_{\tilde{k}=1}^{\mathcal{K}}d_{\tilde{k},t},$ (21)
where $d_{\tilde{k},t},\;t=1,2,\ldots,\mathcal{N}$ is the $t^{th}$ element of $d_{\tilde{k}}=\tilde{\mathcal{W}}_{\tilde{k}}\tilde{\varpi}_{\tilde{k}},\;({\tilde{k}}=1,2,\ldots,\mathcal{K})$. Thus, the time series $Y$ can be expressed as a linear combination of the high-frequency and low-frequency coefficients as
$Y_{t}=S_{\mathcal{K},t}+\sum_{\tilde{k}=1}^{\mathcal{K}}\delta_{\tilde{k},t},$ (22)
where $S_{\mathcal{K},t}=\sum_{j=\tilde{k}+1}^{\mathcal{K}+1}d_{j,t}$ is the $\mathcal{K}^{th}$ level smooth coefficient and $\delta_{\tilde{k},t}=\sum_{j=1}^{\tilde{k}}d_{j,t}$ is the $\tilde{k}^{th}$ level details associated with the MODWT-based MRA decomposition. A salient feature of this scale-based additive transformation is that the wavelet and scaling coefficients of the series are associated with the zero-phase filters. Hence, they are of the same size as the original series and are perfectly aligned with it. A more detailed structural overview of the MODWT process can be found in Percival & Mofjeld (1997).
8.3 Economic filters - Hodrick-Prescott and Christiano-Fitzgerald filters:
The Hodrick-Prescott filter is a popular econometric method that has found profound applications in extracting the business cycle components and as a detrending method for financial time series Hodrick & Prescott (1997). The method is named after its proponents: - Robert J. Hodrick and Edward C. Prescott. This nonparametric method is widely adopted by economists in central banks, international economic agencies, industry, and in government sector Phillips & Shi (2019). There is a wide array of time series filters commonly used in macroeconomic and financial research and analysis to decompose the overall time series into trend, cycle, and irregular components Gençay et al. (2001). Among them, the HP filter has been in vogue among researchers for extracting trend and cycle components from the time series. One key aspect of the HP filter is that it is quite effective and can be applied to non-stationary time series. Therefore, it makes an excellent choice for decomposing CPI inflation, EPU, and GPRC series in this exercise.
For any time series data ($Y_{t}:t=1,\ldots.,\ n$), the HP filter decomposes the series into a trend ($\tau_{t}$), and a cyclical component ($C_{t}$), by solving the following minimization problem with respect to $\tau_{t}$:
${\widehat{\tau}}_{t}^{HP}=\underset{\tau_{t}}{\operatorname{argmin}}\left\{\sum_{t=1}^{n}{({Y_{t}-\tau_{t})}^{2}}+\ \theta\sum_{t=2}^{n}{({\mathrm{\Delta}^{2}\tau_{t})}^{2}}\right\}\;\text{and}\;{\widehat{C}}_{t}^{HP}=Y_{t}-{\widehat{\tau}}_{t}^{HP}$ (23)
[…]
$\mathcal{H}(f,\theta)=\frac{4\theta[1-cos(2\pi f{)]}^{2}}{1+4\theta[1-\cos(2\pi f)]^{2}}$ (24)
and the trend component of the HP filter is given by:
$\tau_{t}=\frac{\phi_{1}\phi_{2}}{\theta}\left[\sum_{j=0}^{\infty}{(A_{1}\phi_{1}^{j}+A_{2}\phi_{2}^{j})Y_{t-j}}+\sum_{j=0}^{\infty}{(A_{1}\phi_{1}^{j}+A_{2}\phi_{2}^{j})Y_{t+j}}\right],$ (25)
where $\phi_{1}\ {and\ \phi}_{2}$ are the complex conjugates and depends on the value of $\theta$and $A_{1}$ and $A_{2}$ are the functions of $\phi_{1}\ and\ \phi_{2}$. By this analysis, it can be argued that the trend component, extracted using the HP filter, is a centered moving average term.
On the other hand, the Christiano-Fitzgerald (CF) (often regarded as a random walk filter) filter is a band pass filter that was primarily constructed using very similar principles as Baxter and King (BK) filter Christiano & Fitzgerald (2003). CF filter has demonstrated better performance compared to the BK filter in the case of real-time applications. CF filter can be calculated as:
$C_{t}=\varphi_{0}Y_{t}+\varphi_{1}Y_{t+1}+\cdots+\varphi_{T-1-t}Y_{T-1}+\tilde{\varphi}_{T-t}Y_{T}+\varphi_{1}Y_{t-1}+\cdots+\varphi_{t-2}Y_{2}+\tilde{\varphi}_{t-1}Y_{1}$ (26)
where $\varphi_{j}=\frac{\sin(jb)-\sin(ja)}{\pi j},j\geq 1$, and $\varphi_{0}=\frac{b-a}{\pi},a=\frac{2\pi}{p_{u}},b=\frac{2\pi}{p_{l}}$ and $\tilde{\varphi}_{k}=-\frac{1}{2}\varphi_{0}-\sum_{j=1}^{k-1}\varphi_{j}$. The parameters, $p_{u}$ and $p_{l}$ represent the break-off point of cycle length in the month. Any cycle value between $p_{u}$ and $p_{l}$ is preserved in the cyclical term $C_{t}$. CF filter aims to formulate the de-trending and smoothing problem in the frequency domain. The CF filter approximates the ideal infinite bandpass filter, which is like the BK filter. One key advantage of CF filters is that the random walk filter can exploit the complete series for the calculation of each filtered point. CF filters can converge well with optimal filters in the long run. This property speaks for the effectiveness of the CF filter as compared to the BK filter in the context of extracting meaningful features (cyclical components) from a specific time series.
In this analysis, the HP filter has been explored to extract the trend component for CPI inflation, EPU, and GPRC series Whereas, the CF filter is used to extract the cyclical terms from all the 3-time series (CPI inflation, EPU, and GPRC) and post that these derived features are used in the downstream FEWNet algorithm to generate the long-term forecasts for the CPI inflation series.
8.4 Proof for Proposition 1
Proof. Using Eq. (7), we write the ARNNx estimator as follows
$\hat{f}=\alpha+\sum_{j^{{}^{\prime}}=1}^{q^{*}}\beta_{j^{{}^{\prime}}}\sigma\left(z_{j^{{}^{\prime}}}\right),$ (27)
where $z_{j^{{}^{\prime}}}=\alpha^{*}+\sum_{i=1}^{r}\beta_{i,j^{{}^{\prime}}}^{*}y_{i}+\gamma^{*}\underline{X}$ where the parameters are estimated by minimizing Eq. (6). For FEWNet, we have
$\begin{gathered}\hat{f}_{0}=\alpha_{0}^{S}+\sum_{i=1}^{q}\beta_{i}^{S}\sigma(\alpha_{i}^{S}+\sum_{j_{0}=1}^{r}{\beta^{S}_{i,j_{0}}}S_{\mathcal{K},j_{0}}+{\gamma^{S}_{i}}^{{}^{\prime}}\underline{X})\\ \hat{f}_{\tilde{k}}=\alpha_{0,\tilde{k}}+\sum_{i=1}^{q}\beta_{i,\tilde{k}}\sigma(\alpha_{i,\tilde{k}}+\sum_{j_{\tilde{k}}=1}^{r}\beta_{i,j_{\tilde{k}}}{\delta}_{\tilde{k},j_{\tilde{k}}}+\gamma^{{}^{\prime}}_{i,\tilde{k}}\underline{X});\;\tilde{k}=1,2,\ldots,\mathcal{K}.\end{gathered}$ (28)
Using Jensen’s inequality and Eq. (2), we get the $\mathcal{R}_{\text{Emp}}$ definition domain is a subspace of $\mathcal{R}_{\text{Emp}}^{W}$
$\operatorname{min}\mathcal{R}_{\text{Emp}}^{W}\leq\operatorname{min}\mathcal{R}_{\text{Emp}}$ (29)
8.5 Conformal Prediction for semi-long-term
Figure 8 depicts the uncertainty quantification of the FEWNet framework for the 12 months ahead forecast horizon. The conformal prediction intervals demonstrated in the figure are of varied widths across countries. In the case of 12 months forecasts for Brazil, the width of the interval averages around 5.06, while for Russia, it hovers around 2.3, for India, it averages 3.7, and the width value for China is around 6.03, the maximum among all the countries. This indicates a higher degree of prediction uncertainty for China than for other countries.
Figure 8: The plot shows the ground truth CPI inflation data (red dots), 12 months point forecasts generated by FEWNet (blue line), AR (green line), XGBoost (purple line), and the conformal prediction interval of FEWNet (yellow shaded) for BRIC Countries